# Taming High-Dimensional Dynamics: Learning Optimal Projections onto Spectral Submanifolds

### v1

@@ IV-C Learning oblique projections from data
5:Return $\mathcal{Y}_{\text{trans}},\dot{\bm{\mathcal{Y}}}_{\text{trans}},\bm{\mathcal{Y}}_{\text{near}},\dot{\bm{\mathcal{Y}}}_{\text{near}},\mathbf{V}_{\text{E}}$
@@
1:$\bm{\mathcal{Y}}_{\text{trans}},\dot{\bm{\mathcal{Y}}}_{\text{trans}},\mathbf{V}_{\text{E}},n_{r}$
@@
$\min_{\mathbf{V}_{\text{opt}},\mathbf{R}_{\text{trans}}}$ $\quad\left\|\mathbf{V}_{\text{opt}}^{\top}\bm{\mathcal{\dot{Y}}}-\mathbf{R}_{\text{trans}}\left(\mathbf{V}_{\text{opt}}^{\top}\bm{\mathcal{Y}}\right)^{1:n_{r}}\right\|^{2}$ (31)
@@
In this section, we learn the oblique projection, reduced dynamics and SSM parametrization from observational data. Observable data comprise of decaying trajectories, which we stack to form a dataset $\bm{\mathcal{Y}}\in\mathbb{R}^{p\times N}$, where $N$ is the trajectory length and $p$ is the dimension of the observable space. Note, $p\geq 2n+1$, to guarantee the embedding of an $n$-dimensional SSM, which commonly is satisfied by enlarging the observable space via delay-embedded trajectories (see [cenedese2022data-driven]).
We partition the dataset into transient $\bm{\mathcal{Y}}_{\text{trans}}\in\mathbb{R}^{p\times N_{T_{1}}}$ and near-manifold $\bm{\mathcal{Y}}_{\text{near}}\in\mathbb{R}^{p\times N-N_{T_{1}}}$ data. Earlier methods ([cenedese2022data-driven]) discarded $\bm{\mathcal{Y}}_{\text{trans}}$, and computed the chart map, reduced dynamics and SSM parametrization in the observable coordinates using $\bm{\mathcal{Y}}_{\text{near}}$. Algorithm 1 recalls these steps and obtains $\mathbf{V}_{\text{E}}$, whose column vectors span the tangent space of the SSM. Setting $\mathbf{V}_{\text{opt}}=\mathbf{V}_{\text{E}}$ recovers an orthogonal projection onto the SSM.
From analysis in Section IV-A, we know $\bm{\mathcal{Y}}_{\text{trans}}$ contains information of local fiberwise directions $\mathbf{V}_{\text{opt}}$. We also leverage the properties of the projection operator (Section IV-B), and devise Algorithm 2 to find an optimal $\mathbf{V}_{\text{opt}}$. We use $(\cdot)^{d_{0}:d}$ to denote the family of all monomials from order $d_{0}$ to $d$. The algorithm employs projected gradient descent or IPOPT [wachter2006implementation] to perform the optimization. Finally, we use the optimized oblique projection $\mathbf{V}_{\text{opt}}$ to learn the reduced dynamics $\mathbf{R}$ and the parameterization map $\mathbf{W}_{\text{nl}}$ with near-manifold data, as detailed in Algorithm 3.
@@
1:$\bm{\mathcal{Y}}_{\text{near}},\dot{\bm{\mathcal{Y}}}_{\text{near}},\bm{\mathcal{Y}}_{u},\dot{\bm{\mathcal{Y}}}_{u},\mathbf{U},\mathbf{V}_{\text{opt}},\mathbf{V}_{\text{E}},n_{r},n_{w}$
@@ V Data-driven control via SSMs
In Section IV-C, we described a method to learn optimal oblique projections for SSM-reduced models using data from uncontrolled systems. To apply this to controlled dynamics, we calibrate the models by applying control inputs $\mathbf{U}$ and collecting the resulting data $\bm{\mathcal{Y}}_{u}$. Using the optimization in Algorithm 3, we compute the control matrix $\mathbf{B}_{r}$. This calibrated model enables closed-loop control with Model Predictive Control (MPC). Unlike standard MPC, our approach uses oblique projections of controlled data with $\mathbf{V}_{\text{opt}}$. For details on the optimal control problem formulation, we refer readers to [alora2025discovering, AloraCenedeseEtAl2023] and Section -D. We now demonstrate the effectiveness of our approach on a benchmark problem and closed-loop control of a continuum trunk robot.  [^5] The source code can be accessed at https://github.com/StanfordASL/Opt-SSM
@@ V-A Slow-Fast System
We consider a two-dimensional system with large timescale separation, adopted from [otto2023learning]. We add generic forcing, $\mathbf{u}=\begin{bmatrix}u_{1}&u_{2}\end{bmatrix}^{\top}$, yielding the governing equations
$\dot{x}_{1}$ $=\lambda x_{1}\left(1-x_{1}^{2}\right)+\alpha u_{1}$ (33)
@@
where $\lambda,\varepsilon>0$ and $\varepsilon^{-1}\gg\lambda$. We set the parameters as $\lambda=\varepsilon=0.1$ and $\alpha=\beta=0.2$. For the uncontrolled system, we collect 10 trajectories around the asymptotically stable fixed point $(1,1)$, and shift the data to be centered around the origin. The chart map and reduced dynamics are found using Algorithm 2 and the parameterization map using Algorithm 3. In Figure 2 (left), we show how SSMs using orthogonal projections compare with those using our optimized oblique projections. We initialize off the slow manifold and, for both methods, project onto it before propagating the reduced dynamics. The orthogonal projection causes the predicted response to be on opposite side of the stable fixed point, resulting in large trajectory error. The oblique projection, on the other hand, projects vertically along the fast dynamics, producing minimal trajectory error.
Recalling the assumed control-affine form of the underlying dynamics (1), we have $\mathbf{B}=\begin{bmatrix}\alpha&\beta\end{bmatrix}^{\top}$, such that the non-autonomous reduced dynamics becomes
$\mathbf{\dot{x}}_{r}=\bm{\mathbf{R}}\mathbf{x}_{r}^{1:n_{r}}+\mathbf{V}_{\text{opt}}^{\top}\mathbf{B}_{r}\mathbf{u}.$ (34)
We simulate the forced response with $u=\begin{bmatrix}e^{-t}&e^{-t}\end{bmatrix}^{\top}$, which causes the system to initially move away from the stable fixed point. The results are shown in Figure 2 (right). Again, orthogonally projecting leads to large errors, while the oblique projection leads to better initialization and captures the amplitude caused by the forcing accurately.
@@ V-B High-Dimensional Trunk Robot
Figure 3: The simulated trunk robot (left) has a 30-link (180 dim.) body actuated by 12 antagonistic tendons ($u\in\mathbb{R}^{6}$). We compare closed-loop tracking performance of the standard SSM scheme with our proposed oblique projection-based SSM.
We collect 10 autonomous decay trajectories for fitting of the chart map, parameterization map and reduced dynamics. Each trajectory is $10s$ in duration, and observations are stored at a time discretization of $dt=0.01s$. We concatenate the observations with 3 time-delays and learn a 5D autonomous SSM model, with $n_{r}=n_{w}=2$. These parameter values align with the findings for similar SSM-reduced models in soft trunk robot geometries (see [kaundinya2025data-driven] for the justification). For the orthogonal projection, we report a mean MSE of $9.16\pm 5.76$ $cm^{2}$ across 10 test decay trajectories, while the oblique projection achieves $1.60\pm 1.10$ $cm^{2}$. An additional 20 controlled trajectories are collected for regression of the control matrix. Specifically, we excite the system by harmonically varying the control inputs.
@@ VI Conclusion and outlook
In this work, we have introduced optimal oblique spectral projections to enhance the accuracy of data-driven model reduction to slow SSMs. By utilizing the stable fiber structure surrounding slow SSMs, we developed a linear approximation of these fibers that reduces long-term prediction errors in the system’s dynamics. Additionally, we have shown how to efficiently learn these projections from data and use them to achieve notable improvements in closed-loop control performance.
Future work could extend our framework by learning nonlinear chart maps, $\mathbf{V}_{\text{opt}}(\mathbf{x}_{r})$, in which case we expect to approximate the curvature of the fibers. Moreover, we assume user-specified transient segmentation, $T_{1}$, which may yield suboptimal results if misspecified. Recent advances in inferring dynamical structures from data [mcinroe2024global] could help automate this partitioning.
@@
Acknowledgements: The authors thank Rohan Sinha for his thoughtful review and Patrick B. Eberhard for assistance with the trunk simulator and visualization.
-A Proof of Lemma 1
Proof:
Note that $\mathbf{w}^{-1}$ exists since $\mathbf{w}$ is a local diffeomorphism in $\mathcal{T}^{\varepsilon}$. Since $\mathbf{w}^{-1}$ is smooth on $\mathcal{T}^{\varepsilon}$ and $\mathcal{T}^{\varepsilon}$ can be chosen small enough such that it is guaranteed to be a compact set (e.g., by shrinking $\epsilon$), then the Jacobian $\mathbf{D}\mathbf{w}^{-1}$ is bounded on $\mathcal{T}^{\varepsilon}$. Thus,
$\left\|\mathbf{D}\mathbf{w}^{-1}(\mathbf{x})\right\|\leq\sigma_{\text{max}}(\mathbf{D}\mathbf{w}^{-1})=\frac{1}{\sigma_{\text{min}}(\mathbf{D}\mathbf{w})}.$
By the Mean Value Theorem, we have that
$\left\|\begin{bmatrix}\mathbf{x}_{r}-\mathbf{x}_{r}^{\prime}\\ \mathbf{x}_{n}-\mathbf{x}_{n}^{\prime}\end{bmatrix}\right\|\leq\frac{1}{\sigma_{\text{min}}(\mathbf{D}\mathbf{w})}\left\|\mathbf{w}(\mathbf{x}_{r},\mathbf{x}_{n})-\mathbf{w}(\mathbf{x}_{r}^{\prime},\mathbf{x}_{n}^{\prime})\right\|,$
which by rearranging, gives us the lower bound. Finally, since $\mathbf{w}$ is differentiable on a convex subset of $(\bm{\sigma}\times\bm{\tau})\big{(}\mathbf{N}^{s,\varepsilon}\big{)}$, then we can apply the mean value theorem such that
$\left\|\mathbf{w}(\mathbf{x}_{r},\mathbf{x}_{n})-\mathbf{w}(\mathbf{x}_{r}^{\prime},\mathbf{x}_{n}^{\prime})\right\|$ $\leq\left\|\mathbf{D}\mathbf{w}\right\|\left\|\begin{bmatrix}\mathbf{x}_{r}-\mathbf{x}_{r}^{\prime}\\ \mathbf{x}_{n}-\mathbf{x}_{n}^{\prime}\end{bmatrix}\right\|$ (35)
$\leq\sigma_{\text{max}}(\mathbf{D}\mathbf{w})\left\|\begin{bmatrix}\mathbf{x}_{r}-\mathbf{x}_{r}^{\prime}\\ \mathbf{x}_{n}-\mathbf{x}_{n}^{\prime}\end{bmatrix}\right\|,$
which gives us the upper bound and completes the proof. ∎
-B Proof of Corollary 1
Proof:
We begin by telescoping the denominator,
$\lim_{t\to\infty}\frac{\left\|\mathbf{F}^{t}(\mathbf{x})-\mathbf{F}^{t}(\mathbf{p})\right\|}{\left\|\mathbf{F}^{t}(\mathbf{x})-\mathbf{F}^{t}(\mathbf{p}^{\prime})\right\|}$ $=\lim_{t\to\infty}\frac{\left\|\mathbf{F}^{t}(\mathbf{x})-\mathbf{F}^{t}(\mathbf{p}^{\prime})\right\|}{\left\|\mathbf{F}^{t}(\mathbf{x})-\mathbf{F}^{t}(\mathbf{p})+\mathbf{F}^{t}(\mathbf{p})-\mathbf{F}^{t}(\mathbf{p})\right\|}$
$\leq\lim_{t\to\infty}\frac{1}{\Big{|}\frac{\left\|\mathbf{F}^{t}(\mathbf{x})-\mathbf{F}^{t}(\mathbf{p})\right\|}{\left\|\mathbf{F}^{t}(\mathbf{x})-\mathbf{F}^{t}(\mathbf{p})\right\|}-\frac{\left\|\mathbf{F}^{t}(\mathbf{p})-\mathbf{F}^{t}(\mathbf{p}^{\prime})\right\|}{\left\|\mathbf{F}^{t}(\mathbf{x})-\mathbf{F}^{t}(\mathbf{p}^{\prime})\right\|}\Big{|}}$
$=\lim_{t\to\infty}\frac{1}{\Big{|}1-\frac{\left\|\mathbf{F}^{t}(\mathbf{p})-\mathbf{F}^{t}(\mathbf{p}^{\prime})\right\|}{\left\|\mathbf{F}^{t}(\mathbf{x})-\mathbf{F}^{t}(\mathbf{p})\right\|}\Big{|}},$
where the second inequality is due to the triangle inequality. By the invariant foliation theorem (Theorem 1(iii)) and since $\mathbf{p}^{\prime}\in\mathcal{F}(\mathbf{p}^{\prime})$, we have that the ratio $\frac{\left\|\mathbf{F}^{t}(\mathbf{p}^{\prime})-\mathbf{F}^{t}(\mathbf{p})\right\|}{\left\|\mathbf{F}^{t}(\mathbf{x})-\mathbf{F}^{t}(\mathbf{p})\right\|}\to\infty$ as $t\to\infty$. Hence,
$\lim_{t\to\infty}\frac{\left\|\mathbf{F}^{t}(\mathbf{x})-\mathbf{F}^{t}(\mathbf{p})\right\|}{\left\|\mathbf{F}^{t}(\mathbf{x})-\mathbf{F}^{t}(\mathbf{p}^{\prime})\right\|}\leq\lim_{t\to\infty}\frac{1}{\Big{|}1-\frac{\left\|\mathbf{F}^{t}(\mathbf{p})-\mathbf{F}^{t}(\mathbf{p}^{\prime})\right\|}{\left\|\mathbf{F}^{t}(\mathbf{x})-\mathbf{F}^{t}(\mathbf{p})\right\|}\Big{|}}=0.$
-C Proof of Lemma 2
Proof:
Since $\mathbf{x}\in\mathcal{F}(\mathbf{p})$, we can denote the reduced coordinates as a function of the normal coordinates, i.e., $\bar{\mathbf{x}}_{r}=\bm{\sigma}(\mathbf{p}),\mathbf{x}_{r}^{\prime}=\bm{\sigma}(\mathbf{p}^{\prime})$. The constant $\gamma_{2}>0$ is chosen such that, for any $\mathbf{p}^{\prime}$ satisfying $\left\|\mathbf{p}-\mathbf{p}^{\prime}\right\|<\gamma_{2}$, we have that $\left\|\mathbf{x}_{r}^{\prime}-\bar{\mathbf{x}}_{r}\right\|^{2}<\eta_{1}\left\|\mathbf{x}_{r}^{\prime}-\bar{\mathbf{x}}_{r}\right\|$ for an appropriately chosen $\eta_{1}>0$.
We derive the upper bound on the flow in the reduced coordinates by Taylor expanding $\varphi_{t}$ at $\mathbf{x}_{r}^{\prime}$,
$\varphi_{t}(\bar{\mathbf{x}}_{r},\mathbf{0})=\varphi_{t}(\mathbf{x}_{r}^{\prime},\mathbf{0})$ $+\mathbf{D}_{1}\varphi_{t}(\mathbf{x}_{r}^{\prime},\mathbf{0})(\mathbf{x}_{r}^{\prime}-\bar{\mathbf{x}}_{r})$ (36)
$+\mathcal{O}(\left\|\mathbf{x}_{r}^{\prime}-\bar{\mathbf{x}}_{r}\right\|^{2}),$
where we used the fact that $\mathbf{D}_{2}\varphi_{t}(\mathbf{x}_{r}^{\prime},\mathbf{0})=\mathbf{0}$. Using this expansion, we construct the following bound
$\left\|\varphi_{t}(\bar{\mathbf{x}}_{r},\mathbf{0})-\varphi_{t}(\mathbf{x}_{r}^{\prime},\mathbf{0})\right\|$ (37)
$\hskip 18.00003pt=\left\|\big{(}\mathbf{D}_{1}\varphi_{t}(\mathbf{x}_{r}^{\prime},\mathbf{0})(\mathbf{x}_{r}^{\prime}-\bar{\mathbf{x}}_{r})+\mathcal{O}(\left\|\mathbf{x}_{r}^{\prime}-\bar{\mathbf{x}}_{r}\right\|^{2})\right\|$
$\hskip 18.00003pt\leq\left\|\mathbf{D}_{1}\varphi_{t}(\mathbf{x}_{r}^{\prime},\mathbf{0})\right\|\left\|\mathbf{x}_{r}^{\prime}-\bar{\mathbf{x}}_{r}\right\|+C_{1}e^{-\mu t}\left\|\mathbf{x}_{r}^{\prime}-\bar{\mathbf{x}}_{r}\right\|^{2}$
$\hskip 18.00003pt\leq(C+C_{1}\eta_{1})e^{-\mu t}\left\|\mathbf{x}_{r}^{\prime}-\bar{\mathbf{x}}_{r}\right\|,$
where the first inequality results from the fact that under Lemma C.1 in [eldering2013normally], if $\left\|\mathbf{D}_{1}\varphi_{t}\right\|\leq Ce^{-\mu t}$, then there exists $\bar{C}$ such that $\left\|\mathbf{D}^{k}_{1}\varphi_{t}\right\|\leq\bar{C}e^{-\mu t}$ for $k\leq r$ and the integral remainder form of a Taylor series [marsden1993elementary] (Theorem 9, p. 195) is bounded by
$\mathcal{O}(\left\|\mathbf{x}_{r}^{\prime}-\bar{\mathbf{x}}_{r}\right\|^{2})$ $\leq\frac{1}{2}\left\|\mathbf{D}^{2}_{1}\varphi_{t}\right\|\left\|\mathbf{x}_{r}^{\prime}-\bar{\mathbf{x}}_{r}\right\|^{2}$ (38)
$\leq\frac{1}{2}\bar{C}e^{-\mu t}\eta_{1}\left\|\mathbf{x}_{r}^{\prime}-\bar{\mathbf{x}}_{r}\right\|.$
In Equation (37), we let $C_{1}=\frac{1}{2}\bar{C}$. Notice that we used Definition 1(iii), and the fact that $\gamma_{2}$ is small enough to ensure $\left\|\mathbf{x}_{r}^{\prime}-\bar{\mathbf{x}}_{r}\right\|^{2}<\eta_{1}\left\|\mathbf{x}_{r}^{\prime}-\bar{\mathbf{x}}_{r}\right\|$. The lemma follows with $K=C+C_{1}\eta_{1}$. ∎
-D Reduced Order Model Predictive Control
This section describes how we implement trajectory tracking using MPC with our learned dynamics model. We formulate the MPC optimization problem as follows
$\underset{\mathbf{u}(\cdot)}{\operatorname{min}}$ $\left\|\delta\mathbf{z}\left(t_{f}\right)\right\|_{\mathbf{Q}_{f}}^{2}+\int_{t_{0}}^{t_{f}}\left(\|\delta\mathbf{z}(t)\|_{\mathbf{Q}}^{2}+\|\mathbf{u}(t)\|_{\mathbf{R}}^{2}\right.$ (39)
$+\left.\|\mathbf{u}(t)-\mathbf{u}(t-\Delta t)\|_{\mathbf{R}_{\Delta}}^{2}\right)dt$
s.t. $\mathbf{x}_{r}(0)=\mathbf{V}_{\text{opt}}^{\top}\left(\mathbf{y}(0)-\mathbf{y}_{\mathrm{eq}}\right)$
$\dot{\mathbf{x}}_{r}(t)=\mathbf{R}\mathbf{x}_{r}(t)^{1:n_{r}}+\mathbf{B}_{r}\mathbf{u}(t)$
$\mathbf{z}(t)=\mathbf{C}\left(\mathbf{V}_{\mathrm{PCA}}\mathbf{x_{r}}(t)+\mathbf{W}_{\mathrm{nl}}\mathbf{x}_{r}(t)^{2:n_{w}}\right)+\mathbf{z}_{\mathrm{eq}},$
$\mathbf{z}(t)\in\mathcal{Z},\quad\mathbf{u}(t)\in\mathcal{U},$
where $\mathbf{z}\in\mathbb{R}^{o}$, with $o\leq p$, is the performance state that we aim to control. Equilibrium quantities are denoted by the subscript ’eq’, as in $y_{\mathrm{eq}}$ and $z_{\mathrm{eq}}$. To solve (39), we discretize the continuous-time dynamics and apply GuSTO [BonalliCauligiEtAl2019]. Note that running this optimization is intended to be real-time feasible, as $n\ll n_{f}$.
[…]

### v2

@@ IV-C Learning Oblique Projections from Data
5:Return $\bm{\mathcal{Y}}_{\text{trans}},\dot{\bm{\mathcal{Y}}}_{\text{trans}},\bm{\mathcal{Y}}_{\text{near}},\dot{\bm{\mathcal{Y}}}_{\text{near}},\mathbf{V}_{\text{E}}$
@@
[…]
@@
$\min_{\mathbf{V}_{\text{opt}},\mathbf{R}_{\text{trans}}}$ $\quad\left\|\mathbf{V}_{\text{opt}}^{\top}\bm{\mathcal{\dot{Y}}}_{\text{trans}}-\mathbf{R}_{\text{trans}}\left(\mathbf{V}_{\text{opt}}^{\top}\bm{\mathcal{Y}}_{\text{trans}}\right)^{1:n_{r}}\right\|^{2}$ (28)
@@
2:Solve the following optimization problems consecutively: $\min_{\mathbf{R}}$ $\left\|\mathbf{V}_{\text{opt}}^{\top}\dot{\bm{\mathcal{Y}}}_{\text{near}}-\mathbf{R}(\mathbf{V}_{\text{opt}}^{\top}\bm{\mathcal{Y}}_{\text{near}})^{1:n_{r}}\right\|^{2}$ (29) $\min_{\mathbf{W}_{\text{nl}}}$ $\left\|\bm{\mathcal{Y}}_{\text{near}}-\mathbf{V}_{\text{E}}(\mathbf{V}_{\text{opt}}^{\top}\bm{\mathcal{Y}}_{\text{near}})-\mathbf{W}_{\text{nl}}(\mathbf{V}_{\text{opt}}^{\top}\bm{\mathcal{Y}}_{\text{near}})^{2:n_{w}}\right\|^{2}$ s.t. $\quad\mathbf{V}_{\text{opt}}^{\top}\mathbf{W}_{\text{nl}}=\mathbf{0}$ $\min_{\mathbf{B}_{r}}$ $\left\|\mathbf{V}_{\text{opt}}^{\top}\dot{\bm{\mathcal{Y}}}_{u}-\mathbf{R}(\mathbf{V}_{\text{opt}}^{\top}\bm{\mathcal{Y}}_{u})^{1:n_{r}}-\mathbf{B}_{r}\mathbf{U}\right\|^{2}$
@@
For real-world systems, we must learn the oblique projection, $\Pi$, from observational data. Specifically, we fit the chart and parameterization maps as defined in (21) and the on-manifold dynamics, $\mathbf{r}:\mathbb{R}^{n}\to\mathbb{R}^{n}$. The chart and parameterization maps must satisfy the invertibility relations [3], $\mathbf{y}=(\bm{\omega}\circ\bm{\nu})(\mathbf{y})$ and $\mathbf{x}_{r}=(\bm{\nu}\circ\bm{\omega})(\mathbf{x}_{r})$, where we replaced $\mathbf{x}$ with the observed states, $\mathbf{y}\in\mathbb{R}^{p}$. The exact form satisfying these relations we choose is
$\mathbf{x}_{r}$ $=\bm{\nu}(\mathbf{y})=\mathbf{V}_{\text{opt}}^{\top}\mathbf{y},$ (30)
$\mathbf{y}$ $=\bm{\omega}(\mathbf{x}_{r})=\mathbf{V}_{\text{E}}\mathbf{x}_{r}+\mathbf{W}_{\text{nl}}\mathbf{x}_{r}^{2:n_{w}},$
since $\mathbf{V}_{\text{opt}}^{\top}\mathbf{V}_{\text{E}}={\mathbf{I}}$ and we will enforce $\mathbf{V}_{\text{opt}}^{\top}\mathbf{W}_{\text{nl}}=\mathbf{0}$. We use $\mathbf{x}_{r}^{2:n_{w}}$ to denote the family of all monomials from order 2 to $n_{w}$. The dynamics on $\mathcal{M}$ takes a similar polynomial form
$\dot{\mathbf{x}}_{r}=\mathbf{r}(\mathbf{x}_{r})\vcentcolon=\mathbf{R}\mathbf{x}_{r}^{1:n_{r}},$ (31)
with $n_{r}$ being the desired order of the Taylor expansion.
To guarantee the embedding of an $n$-dimensional SSM, we require $p\geq 2n+1$, which is commonly satisfied by enlarging the observable space via delay-embedded trajectories (see [2]). Differential embeddings are locally observable [24], which implies that delay embeddings are also likely locally observable (since the latter is a discrete version of the former). Thus, if the robot’s sensors are appropriately placed to detect the dominant modes [25] and the robot’s controlled behaviors keep the system close to the SSM, the reduced state in (31) is approximately observable.
Observable data comprise of decaying trajectories, which we stack to form a dataset $\bm{\mathcal{Y}}\in\mathbb{R}^{p\times N}$, where $N$ is the trajectory length and $p$ is the dimension of the observable space. We partition this dataset into transient and near-manifold components: $\bm{\mathcal{Y}}_{\text{trans}}\in\mathbb{R}^{p\times N_{T_{1}}}$ contains the first $N_{T_{1}}$ time steps corresponding to the transient dynamics, and $\bm{\mathcal{Y}}_{\text{near}}\in\mathbb{R}^{p\times(N-N_{T_{1}})}$ captures the remaining near-manifold data. Earlier methods, for example [2], discarded $\bm{\mathcal{Y}}_{\text{trans}}$, and computed the chart map, reduced dynamics and SSM parametrization in the observable coordinates using $\bm{\mathcal{Y}}_{\text{near}}$. Algorithm 1 recalls these steps and obtains $\mathbf{V}_{\text{E}}$, whose column vectors span the tangent space of the SSM. Setting $\mathbf{V}_{\text{opt}}=\mathbf{V}_{\text{E}}$ recovers an orthogonal projection onto the SSM.
From analysis in Section IV-A, we know $\bm{\mathcal{Y}}_{\text{trans}}$ contains information of local fiberwise directions $\mathbf{V}_{\text{opt}}$. Algorithm 2 therefore optimizes jointly over the dynamics in (31) and the projection direction to find an optimal $\mathbf{V}_{\text{opt}}$. Note that the constraint in (28) enforces that the projection operator is idempotent (recall Theorem 3(i)). We employ projected gradient descent or IPOPT [26] to solve this non-convex optimization problem. Finally, we use the optimized oblique projection $\mathbf{V}_{\text{opt}}$ and near-manifold data to learn the coefficient matrices of the reduced dynamics and the parameterization map, $\mathbf{R}$ and $\mathbf{W}_{\text{nl}}$, as detailed in Algorithm 3. Indeed, the invertibility relations are satisfied as a result of both the objective function and the constraint in the optimization of $\mathbf{W}_{\text{nl}}$.
It is important to note that learning $\Pi$ from data far from the fixed point minimizes average error by fitting prevalent transients. While recovering the true nonlinear fibers would exactly minimize (15), this typically requires prohibitive amounts of data. Our method therefore strikes a balance between accuracy and practicality.
@@ V Data-Driven Control via SSMs
In Section IV-C, we described a method to learn optimal oblique projections for SSM-reduced models using data from uncontrolled systems. To apply this to controlled dynamics, we augment the autonomous reduced dynamics with a linear control contribution, i.e., with slight abuse of notation,
$\dot{\mathbf{x}}_{r}=\mathbf{r}(\mathbf{x}_{r})+\mathbf{B}_{r}\mathbf{u}.$
We apply control inputs to excite the system and collect the resulting observables, yielding the stacked matrices $\mathbf{U}\in\mathbb{R}^{m\times N^{\prime}}$ and $\bm{\mathcal{Y}}_{u}\in\mathbb{R}^{p\times N^{\prime}}$. Using the optimization in Algorithm 3, we compute the control matrix $\mathbf{B}_{r}$. This calibrated model enables closed-loop control with Model Predictive Control (MPC). Unlike standard MPC, our approach uses oblique projections of controlled data with $\mathbf{V}_{\text{opt}}$. For details on the optimal control problem formulation, we refer readers to Section -E. We now demonstrate the effectiveness of our approach on a benchmark problem and closed-loop control of a continuum trunk robot.  [^4] All necessary code to reproduce the results can be accessed at https://github.com/StanfordASL/Opt-SSM.
@@ V-A Slow-Fast System
To illustrate the importance of projection direction, we consider a two-dimensional system with large timescale separation, adopted from [15]. We add forcing, $\mathbf{u}=\begin{bmatrix}u_{1}&u_{2}\end{bmatrix}^{\top}$, such that the governing equations become
$\dot{x}_{1}$ $=\lambda x_{1}\left(1-x_{1}^{2}\right)+\alpha u_{1},$ (32)
@@
where $\lambda,\varepsilon>0$ and $\varepsilon^{-1}\gg\lambda$. We set the parameters as $\lambda=\varepsilon=0.1$ and $\alpha=\beta=0.2$. For the uncontrolled system, we collect 10 trajectories around the asymptotically stable fixed point $(1,1)$, and shift the data to be centered around the origin. In this example, we observe the full state space and reduce to a 1-dimensional system.
The chart map and reduced dynamics are found using Algorithm 2 and the parameterization map using Algorithm 3. In Figure 2 (left), we show how SSMs using orthogonal projections compare with those using our optimized oblique projections. We initialize off the slow manifold and, for both methods, project onto it before propagating the reduced dynamics. The orthogonal projection causes the predicted response to be on opposite side of the stable fixed point, resulting in large trajectory error. The oblique projection, on the other hand, projects vertically along the fast dynamics, producing minimal trajectory error. Additionally, we simulate the forced response with $\mathbf{u}=\begin{bmatrix}e^{-t}&e^{-t}\end{bmatrix}^{\top}$. The results in Figure 2 (right) show again improved prediction accuracy.
Figure 3: The simulated trunk robot (left) has a 30-link (180 dim.) body actuated by 12 antagonistic tendons ($\mathbf{u}\in\mathbb{R}^{6}$). We compare closed-loop tracking performance of the standard SSM scheme with our proposed oblique projection-based SSM.
@@ V-B High-Dimensional Trunk Robot
We collect 10 autonomous decay trajectories for fitting of the chart map, parameterization map and reduced dynamics. Each trajectory is $10s$ in duration, and observations are stored at a time discretization of $dt=0.01s$. We concatenate the observations with 3 time-delays and learn a 5D autonomous SSM model, with $n_{r}=n_{w}=2$. These parameter values align with the findings for similar SSM-reduced models in soft trunk robot geometries (see [6] for the justification). For the orthogonal projection, we report a mean MSE of $9.16\pm 5.76$ $cm^{2}$ across 10 test decay trajectories, while the oblique projection achieves $1.60\pm 1.10$ $cm^{2}$. An additional 20 controlled trajectories are collected for regression of the control matrix. Specifically, we excite the system by harmonically varying the control inputs.
@@ VI Conclusion and Outlook
In this work, we have introduced optimal oblique spectral projections to enhance the accuracy of data-driven model reduction to slow SSMs. By utilizing the stable fiber structure surrounding slow SSMs, we developed a linear approximation of these fibers that reduces long-term prediction errors in the system’s dynamics. Additionally, we have shown how to effectively learn these projections from data and use them to achieve notable improvements in closed-loop control performance.
Future work could extend our framework by learning nonlinear chart maps, $\mathbf{V}_{\text{opt}}(\mathbf{x}_{r})$, in which case we expect to approximate the curvature of the fibers. Moreover, we assume user-specified transient segmentation, $T_{1}$, which may yield suboptimal results if misspecified. Recent advances in inferring dynamical structures from data [27] could help automate this partitioning.